OpenStudy (anonymous):

F (×)=(1/2•[x] D:{x|xeR;-1

OpenStudy (zzr0ck3r):

what is the question?

OpenStudy (anonymous):

It says draw the graph of the step function for the following domain be sure to label number both axis

OpenStudy (anonymous):

I just need to know how to do it step by step plz

OpenStudy (zzr0ck3r):

What do you get for \(x=1\)?

OpenStudy (zzr0ck3r):

Should that be \(f(x) = (\frac{1}{2})\lceil x \rceil\)?

OpenStudy (anonymous):

Yes that's what's on there

OpenStudy (zzr0ck3r):

So what this say, is take a number, find the next biggest integer and cut that in half Example f(2)=1 f(1/2) = 1/2 f(1.2) = 1

OpenStudy (anonymous):

Right but I don't know how to graph ot

OpenStudy (anonymous):

It

OpenStudy (zzr0ck3r):

what will any number between 0 and 1 evaluate to?

OpenStudy (anonymous):

0

OpenStudy (zzr0ck3r):

take any integer between -1 and 0, the next biggest integer is 0, cut that in half to get 0 take any number between 0 and 1, the next biggest integer is 1, cut that in half and you get 1/2 take any number between 1 and 2, the next biggest integer is 2, cut that in half to get 1. Take any number between 2 and 3, the next biggest integer is 3, cut that in half and get 1.5

OpenStudy (anonymous):

Thanks

OpenStudy (zzr0ck3r):

also, now that you understand how this works. http://www.wolframalpha.com/input/?i=%281%2F2%29ceileng%28x%29